OpenStudy (eyka94):

Hello! Help me solve this please : lim x→1 (1/x - 1)/(x-1). The answer is -1. I'd like to know the working. :(

OpenStudy (unklerhaukus):

\[\lim\limits_{ x→1 }\left(\frac{\frac 1{x}-1}{x-1}\right)\] ?

OpenStudy (unklerhaukus):

or\[\lim\limits_{ x→1 }\left(\frac{\frac 1{x-1}}{x-1}\right)\] ?

OpenStudy (eyka94):

UNCLERHAUKUS, the first one that you posted.

OpenStudy (anonymous):

are you familiar wit L'hopital rule ??

OpenStudy (eyka94):

@sami-21 I've learned that (L'hospital rule) in college but I hardly understand it. Lol. Is it easier?

OpenStudy (eyka94):

@UnkleRhaukus Yeahhh the first one that you posted. (I couldn't copy the equation.)

OpenStudy (anonymous):

I think it is very simple...

OpenStudy (eyka94):

@waterineyes Help me please. I suck at Calculus :|

OpenStudy (anonymous):

it is Just take derivative of numerator and denominator then apply the limits. \[\Large \frac{\frac{d}{dx}(\frac{1}{x}-1)}{\frac{d}{dx}(x-1)}\] after taking derivative we are left with \[\Large \frac{-1}{x^2}\] (deliminator gets one ) now apply the limits \[\Large \lim_{x \rightarrow 1}(\frac{-1}{x^2})=\frac{-1}{1}=-1\] i hope it is clear

OpenStudy (anonymous):

@eyka94 is it ok now ?

OpenStudy (anonymous):

\[\lim\limits_{ x→1 }\left(\frac{\frac 1{x}-1}{x-1}\right) \implies \lim_{x \rightarrow 1}(\frac{1-x}{x(x-1)}) \implies (-1)\lim_{x \rightarrow 1}\frac{x-1}{x(x-1)}\] \[\implies (-1) \lim_{x \rightarrow 1}\frac{1}{x}\] Plug in the limits now..

OpenStudy (eyka94):

@sami-21 Thank you. But this question is under "Limits and Continuity" chapter. Can we answer it in a "LIMIT" way. :P

OpenStudy (anonymous):

There is no need of derivative..

OpenStudy (anonymous):

yes you can do it.

OpenStudy (eyka94):

@waterineyes @sami-21 Thanks to you both! <3 Well it's my first day here and barely know anything about how to use this website haha. Thanks again. ;)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

you 're welcome :)

OpenStudy (anonymous):

\[\huge \color{blue}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (eyka94):

THANK I LOVE YOU ALL!